Two point charges q1 = 25nC and q2 = -75nC are separated by a distance of 3.0 cm. Find the magnitude and direction of (a) the electric force that q1 exerts on q2 and (b) the electric force that q2 exerts on q1.


1
Expert's answer
2022-03-04T08:31:13-0500
"F=k\frac{q_1q_2}{r^2}""F=9*10^9*\frac{25*10^{-9}*75*10^{-9}}{(0.03)^2}=0.019\:\rm N"


(a) toward q1

(b) toward q2